# The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards

### 11 Apr 2026 v2

@@ 4.2 Metric Fragility and LLM Judges
While RLVR often targets verifiable domains (math/coding), many evaluation targets either lack programmatic verifiers or have only partial/weak ones (e.g., safety/refusal appropriateness, long-form coherence and rationale quality, multi-turn task success, or fuzzy information extraction). The evaluation often relies on LLM judges. LLM-judge pipelines offer convenience but can be brittle: small changes (seed choice, $k$, dataset versioning, instruction placement, option order, or even tensor-parallel settings) produce swings comparable to reported gains (Sun et al., 2025a). Judges are also manipulable (Zhao et al., 2025). Beyond judges, seemingly modest metric or setup shifts can alter conclusions about “stable reasoning” (Liu et al., 2024a). Where programmatic verifiers exist, they should be preferred. When judges are unavoidable, prompt/format perturbations and adversarial probes should be part of the protocol, with robustness deltas reported, inter-judge agreement documented, and all templates/configurations released to reduce hidden degrees of freedom.
@@ 4.3 Calibration Drift and Overconfidence
AceReason-Nemotron-14B AIME-24 78.60 (avg@64; Yang et al. 2025) 77.29 +1.31+1.31
AIME-25 67.40 (avg@64; Yang et al. 2025) 66.04 +1.36+1.36
DAPO-Qwen-32B AIME-24 50.00 (accuracy; Yu et al. 2025) 51.56 1.56-1.56
@@
STILL-3-1.5B AIME-24 39.33 (avg@64; T=0.6T{=}0.6, top-p=0.95p{=}0.95; Min et al. 2025) 31.46 +7.87+7.87
@@
Polaris-7B-Preview AIME-24 72.60 (avg@32; NLP and collaborators 2025) 66.46 +6.14+6.14
AMC-23 89.00 (avg@8; NLP and collaborators 2025) 93.59 4.59-4.59
@@
To make these issues concrete, we compared widely cited checkpoints to budget parity controlled runs using the same verifier, matched decoding budgets, and aligned dataset versions.
Important: some sources report pass@1 while others report avg@64 or alternative aggregation (e.g., majority vote) under different decoding settings; in those cases, $\Delta$ conflates metric and budget differences with genuine policy differences. We therefore treat Table 2 as evidence of evaluation sensitivity. Furthermore, the gaps in Table 2 are exacerbated by three main factors:
@@
Multi-sample reporting (e.g., avg@64 with T=0.6T{=}0.6, top-p=0.95p{=}0.95 for STILL-3-1.5B) inflates scores relative to single-shot pass@1 under the same verifier (Min et al., 2025). In our standardized runs, this manifests as +7.87+7.87 for STILL and +6.14+6.14 for Polaris-7B on AIME-24 (NLP and collaborators, 2025).
@@
Changes in prompt templates and dataset slices (e.g., AIME curation, AMC answer formats) shift accuracy by several points, with exact hashes frequently missing from model cards (NLP and collaborators, 2025; Liu and contributors, 2025). This helps explain mixed signs for Nemotron-1.5B v1 (Minerva +8.71+8.71 vs. AMC-23 6.41-6.41).
@@
Sources sometimes mix “accuracy,” pass@k, avg@k (mean of pass@1 across draws), and maj@k (majority vote), or substitute judges where verifiers exist. Once we normalize to a single verifier/spec, several deltas compress (e.g., DAPO-32B on AIME-24: 1.56-1.56(Sun et al., 2025a; Zhao et al., 2025).
@@ 5 Perspective III: Data Contamination
Data provenance is a first-order confound: if pretraining or RL data overlap evaluation sets, measured “reasoning” may reflect memorization. Contamination-aware evaluations in coding (Liu et al., 2024b) and multi-domain leaderboards (Liang et al., 2023) already recommend strict provenance checks; our partial-prompt audit provides direct evidence in math. Partial-prompt reconstruction is a high-precision but not exhaustive signal of overlap: we treat it as one useful probe to be triangulated with lexical and fuzzy matching and with fresh test sets.
@@
Following Wu and others (2025), we reveal only the first $x\%$ of each problem ($x\in\{80,60,40\}$), greedily decode the suffix, and compute ROUGE-L@$x$, EM@$x$, and ACC@$x$. High EM/ACC under large prefixes indicates the model can reconstruct the hidden tail and final answer verbatim.
@@
As summarized in Table 3, Qwen2.5-Math-7B and Qwen3-14B-Base achieve high ACC@80 on legacy math sets (around 58–60% on MATH-500) yet collapse on AIME-2025 (0–3%). Llama-3-1.8B remains near zero across the same sets, consistent with clean evaluation. This pattern (strong on older math sets, absent on the newest release) is consistent with substantial overlap between legacy math benchmarks and the training corpora of some Qwen-family models, and with comparatively cleaner evaluation for Llama. (See Appendix Tables 69 for ACC/ROUGE/EM at 80/60/40% prefixes.) Similar contamination-aware stance and methodology are advocated in code by Liu et al. (2024b) and in broader evaluations by Liang et al. (2023).
@@ 6 The Tax-Aware Minimum Standard for RLVR
Under the minimum standard, each reported accuracy figure must be accompanied by: (i) refusal or “not attempted” rates, (ii) shared accuracy on the intersection of items both models attempted, and (iii) a calibration metric such as ECE computed from the model’s confidence scores (Hochlehnert et al., 2025; Leng et al., 2024). For settings that rely on LLM-as-a-judge rather than programmatic verifiers, we also require at least one judge stress test: the same outputs scored under several prompt templates or instruction orderings, with the spread in scores reported (Zhao et al., 2025; Sun et al., 2025a; Yang et al., 2026). This does not remove judge fragility, but it makes visible how sensitive a claimed RLVR gain is to small changes in the judge pipeline. Together, these measurements ensure that gains are not driven primarily by attempt inflation, miscalibration, or a brittle judge configuration.
@@
Third, RLVR claims must be supported by explicit data provenance checks. section 5 used partial-prompt completion to show that some math benchmarks are heavily memorized by Qwen families: the model can reconstruct the hidden tail and final answer when given 80 percent of the problem on legacy sets, but not on fresh ones such as AIME-2025. Without such checks, apparent reasoning gains can simply reflect better recall of training data (Liu et al., 2024b; Liang et al., 2023; Wu and others, 2025).
@@ 7 Conclusion
Broader relevance. The same pressures arise in SFT, RLHF, and test-time compute (e.g., CoT and maj@k). This tax-aware minimum standard applies unchanged beyond RLVR.
Why focus on RLVR. We focus on RLVR because verifiers make these trade-offs measurable and make RLVR a useful proving ground. As a result, it is a useful proving ground for reliability methods that transfer to broader LLM training and deployment. We hope this synthesis helps the community separate genuine capability expansion from artifacts of budgets, metrics, and data, and encourages tax-aware training and reporting across RLVR, SFT, and RLHF.
@@
Acknowledgments
This work was supported in part by Google Cloud credits through the Gemini Academic Program, which enabled large-scale evaluation and experimentation. We also thank DSO National Laboratories (Singapore) for support on knowledge-focused reasoning research. The views and conclusions expressed in this paper are those of the authors and do not necessarily reflect the official policies of the sponsors.
@@ A.1 Supplementary Figure for the Introduction
This appendix subsection contains Figure 2, which shows the RLVR activity trend alongside AIME performance.
@@ A.2 Potential Criticisms and Responses
Data quality and curricula matter (Chen et al., 2025a; Wen et al., 2025a), but reward/metric design independently shapes failure modes (determinacy/overconfidence under judge rewards; terse/off-topic outputs under naive factuality; calibration drift) (Chen et al., 2025b; Leng et al., 2024; Hochlehnert et al., 2025). We treat both data and objectives as first-class knobs (see section 6).
@@
When verifiers are unavailable or partial, we use LLM judges with robustness probes (prompt/format perturbations, adversarial tests) and publish templates/configs; judges are manipulable, so robustness deltas are reported (Zhao et al., 2025; Sun et al., 2025a). Prefer verifiers when possible (section 4).

### 25 May 2026 v3

@@ 4.2 Metric Fragility and LLM Judges
While RLVR often targets verifiable domains (math/coding), many evaluation targets either lack programmatic verifiers or have only partial/weak ones (e.g., safety/refusal appropriateness, long-form coherence and rationale quality, multi-turn task success, or fuzzy information extraction). The evaluation often relies on LLM judges. LLM-judge pipelines offer convenience but can be brittle: small changes (seed choice, $k$, dataset versioning, instruction placement, option order, or even tensor-parallel settings) produce swings comparable to reported gains (Sun et al., 2025a). Judges are also manipulable (Zhao et al., 2025b). Beyond judges, seemingly modest metric or setup shifts can alter conclusions about “stable reasoning” (Liu et al., 2024). Where programmatic verifiers exist, they should be preferred. When judges are unavoidable, prompt/format perturbations and adversarial probes should be part of the protocol, with robustness deltas reported, inter-judge agreement documented, and all templates/configurations released to reduce hidden degrees of freedom.
@@ 4.3 Calibration Drift and Overconfidence
AceReason-Nemotron-14B AIME-24 78.60 (avg@64; Liu et al. 2025c) 77.29 +1.31+1.31
AIME-25 67.40 (avg@64; Liu et al. 2025c) 66.04 +1.36+1.36
DAPO-Qwen-32B AIME-24 50.00 (avg@32; T=1.0T{=}1.0, top-p=0.7p{=}0.7; Yu et al. 2025) 51.56 1.56-1.56
@@
STILL-3-1.5B AIME-24 39.33 (pass@1/accuracy; 5 sampled responses; Chen et al. 2025c) 31.46 +7.87+7.87
@@
Polaris-7B-Preview AIME-24 72.60 (avg@32; POLARIS-Project 2025) 66.46 +6.14+6.14
AMC-23 89.00 (avg@8; POLARIS-Project 2025) 93.59 4.59-4.59
@@
To make these issues concrete, we compared widely cited checkpoints to budget-parity controlled runs using the same verifier, matched decoding budgets, and aligned dataset versions.
Important: some sources report pass@1 while others report avg@k, sampled accuracy, or alternative aggregation under different decoding settings; in those cases, $\Delta$ conflates metric and budget differences with genuine policy differences. We therefore treat Table 2 as evidence of evaluation sensitivity. Furthermore, the gaps in Table 2 are exacerbated by three main factors:
@@
Multi-sample reporting and prompt/decoding choices can materially change reported scores. For example, AceReason-Nemotron-14B reports avg@64, DAPO-Qwen-32B reports avg@32, and STILL-3-1.5B reports 39.33 on AIME-24 under its own sampled pass@1 evaluation (Liu et al., 2025c; Yu et al., 2025; Chen et al., 2025c). In our standardized runs, this manifests as +7.87+7.87 for STILL and +6.14+6.14 for Polaris-7B on AIME-24 (POLARIS-Project, 2025).
@@
Changes in prompt templates and dataset slices (e.g., AIME curation, AMC answer formats) shift accuracy by several points, with exact hashes frequently missing from model cards (POLARIS-Project, 2025; TIGER-Lab, 2025). This helps explain mixed signs for Nemotron-1.5B v1 (Minerva +8.71+8.71 vs. AMC-23 6.41-6.41).
@@
Sources sometimes mix “accuracy,” pass@k, avg@k (mean of pass@1 across draws), and maj@k (majority vote), or substitute judges where verifiers exist. Once we normalize to a single verifier/spec, several deltas compress (e.g., DAPO-32B on AIME-24: 1.56-1.56(Sun et al., 2025a; Zhao et al., 2025b).
@@ 5 Perspective III: Data Contamination
Data provenance is a first-order confound: if pretraining or RL data overlap evaluation sets, measured “reasoning” may reflect memorization. Contamination-aware evaluations in coding (Jain et al., 2024) and multi-domain leaderboards (Liang et al., 2023) already recommend strict provenance checks; our partial-prompt audit provides direct evidence in math. Partial-prompt reconstruction is a high-precision but not exhaustive signal of overlap: we treat it as one useful probe to be triangulated with lexical and fuzzy matching and with fresh test sets.
@@
Following Wu et al. (2026b), we reveal only the first $x\%$ of each problem ($x\in\{80,60,40\}$), greedily decode the suffix, and compute ROUGE-L@$x$, EM@$x$, and ACC@$x$. High EM/ACC under large prefixes indicates the model can reconstruct the hidden tail and final answer verbatim.
@@
As summarized in Table 3, Qwen2.5-Math-7B and Qwen3-14B-Base achieve high ACC@80 on legacy math sets (around 58–60% on MATH-500) yet collapse on AIME-2025 (0–3%). Llama-3-1.8B remains near zero across the same sets, consistent with clean evaluation. This pattern (strong on older math sets, absent on the newest release) is consistent with substantial overlap between legacy math benchmarks and the training corpora of some Qwen-family models, and with comparatively cleaner evaluation for Llama. (See Appendix Tables 69 for ACC/ROUGE/EM at 80/60/40% prefixes.) Similar contamination-aware stance and methodology are advocated in code by Jain et al. (2024) and in broader evaluations by Liang et al. (2023).
@@ 6 The Tax-Aware Minimum Standard for RLVR
Under the minimum standard, each reported accuracy figure must be accompanied by: (i) refusal or “not attempted” rates, (ii) shared accuracy on the intersection of items both models attempted, and (iii) a calibration metric such as ECE computed from the model’s confidence scores (Hochlehnert et al., 2025; Leng et al., 2024). For settings that rely on LLM-as-a-judge rather than programmatic verifiers, we also require at least one judge stress test: the same outputs scored under several prompt templates or instruction orderings, with the spread in scores reported (Zhao et al., 2025b; Sun et al., 2025a; Yang et al., 2026). This does not remove judge fragility, but it makes visible how sensitive a claimed RLVR gain is to small changes in the judge pipeline. Together, these measurements ensure that gains are not driven primarily by attempt inflation, miscalibration, or a brittle judge configuration.
@@
Third, RLVR claims must be supported by explicit data provenance checks. section 5 used partial-prompt completion to show that some math benchmarks are heavily memorized by Qwen families: the model can reconstruct the hidden tail and final answer when given 80 percent of the problem on legacy sets, but not on fresh ones such as AIME-2025. Without such checks, apparent reasoning gains can simply reflect better recall of training data (Jain et al., 2024; Liang et al., 2023; Wu et al., 2026b).
@@ A.1 Supplementary Figure for the Introduction
This appendix subsection contains Fig. 2, which shows the RLVR activity trend alongside AIME performance.
@@ A.2 Potential Criticisms and Responses
Data quality and curricula matter (Chen et al., 2026; Wen et al., 2025a), but reward/metric design independently shapes failure modes (determinacy/overconfidence under judge rewards; terse/off-topic outputs under naive factuality; calibration drift) (Chen et al., 2025a; Leng et al., 2024; Hochlehnert et al., 2025). We treat both data and objectives as first-class knobs (see section 6).
@@
When verifiers are unavailable or partial, we use LLM judges with robustness probes (prompt/format perturbations, adversarial tests) and publish templates/configs; judges are manipulable, so robustness deltas are reported (Zhao et al., 2025b; Sun et al., 2025a). Prefer verifiers when possible (section 4).